In a mixer enters two streams: CO2 at a rate of 2kg/s and 35oC as well as N2 at 

a rate of 3 kg/s and 50oC. During mixing, 100kW of heat is added while all inlet 

and outlet streams are at 4 bar. Obtain outlet mixture temperature, partial 

pressures of each gas.


1
Expert's answer
2021-07-07T11:41:57-0400

Question

In a mixer enters two streams :C02 at a rate of 2kg/s and 35°c as well as N2 at a rate of 3kg/s and 50°c.During mixing,100kW of heat is added while all inlet and outlet streams are at 4 bar. Obtain outlet mixture temperature ,partial pressures of each gas.

Solution

Mass balance;

"m_{CO_2}+m_{N_2} =m_m"

Where mm is the mass of mixture

mm=2kg/s +3kg/s=5kg/s

Energy balance;

"Q_{in}+m_{CO_2}h_{CO_2}"+m"_{N_2}h_{N_2}"=m"_mh_m"

Qin is the heat added during mixing.

Since both Nitrogen and Carbon dioxide are diatomic gases, their specific heat capacity at constant pressure is give as follows;

"C_p =\frac 7 2 \times R"

Where R= "\frac {R^~}M"

R=8.314

M is Molar mass for each gas.

"C_p" for Nitrogen;

"R_{N_2}" ="\frac {8.314}{28}"

"C_p=\frac {8.314} {28} \times \frac 7 2" =1.03925kJ/kgK

"C_p" for Carbon dioxide;

"R_{CO_2}=\frac {8.314} {44}"

"C_p =\frac {8.314}{44} \times \frac 7 2"=0.6613kJ/kgK

"C_p" for the mixture;

mm =5kg/s

Mass fraction mf for each gas in mixture:

"mf_{CO_2}" ="\frac 2 5" =0.4

"mf_{N_2}=\frac 3 5" =0.6

Number of moles in mixture "n_m" ;

n="\frac m M"

"n_m =n_{CO_2} +n_{N_2}"

"n_m=\frac 2 {44} +\frac 3{28}" =0.1526

Mole fractions y;

"y_{CO_2}=\frac {0.04545} {0.1526}=0.2979"

"y_{N_2}=\frac {0.1071}{0.1526}=0.7121"

Molar mass of mixture Mm;

"M_m=y_{CO_2}M_{CO_2} +y_{N_2}M_{N_2}"

"M_m=(0.2979\times{44} )+(0.7021\times{28})"

Mm=32.77

Rm="\frac {8.314}{32.77}=0.2537"

Cp="\frac 7 2\times 0.2537" =0.8880kJ/kgK

Use the energy equation to find the Temperature of the mixture;

Since h=CpT

100+2(0.6613"\times"308)+3(1.03925"\times"323)=5×0.8880×Tm

Tm=341.04K or 68.04°C

Partial pressures for individual gases;

P=y×Pm

"P_{CO_2}=0.2979\u00d74=1.1916" bar

"P_{N_2}=0.7021\u00d74=2.8084" bar